OpenStudy (anonymous):

Find the exact value by using a half-angle identity. Sin(7pi/8)

OpenStudy (lynfran):

ok so the half angle identity for sin is.....\[\sin \frac{ x }{ 2 }=\sqrt{(1-cosx)/2}\] so \[\sqrt{(1-\cos7\pi)/8} \]\[\sqrt{(1/4)}\sqrt{(1-\cos7\pi)/2}\] we end up with \[\frac{ 1 }{ 2 }*\sqrt{(1-(-1))/2}\] \[\frac{ 1 }{ 2 }*\sqrt{(1+1)/2}\] \[\frac{ 1 }{ 2 }*\sqrt{1}\] ok u can finish it from there?

OpenStudy (lynfran):

so did u get the answer?

OpenStudy (lynfran):

\[\pm \frac{ 1 }{ 2 }\]